Probability

Get insights from 304 questions on Probability, answered by students, alumni, and experts. You may also ask and answer any question you like about Probability

  • 304

    Questions

  • 0

    Discussions

  • 0

    Active Users

  • 4

    Followers

Have a question about Probability? Ask our experts

Ask Your Question

Keep it short & simple. Type complete word. Avoid abusive language.

New answer posted•a year ago
S
a year ago

This is a  Short Answer Type Questions as classified in NCERT Exemplar

Sol:

Here,  we  have  X=0, 1, 2, 3         [?  die  is  thrown  3  times]and  p=16,  q=1−16=56∴  P(X=0)=P(not 2).P(not 2).P(not 2)=56.56.56=125216      P(X=1)=P(2).P(2).P(not 2)+P(2).P(not 2).P(2)+P(not 2).P(2).P(2)                            =16.16.56+16.56.16+56.16.16=5216+5216+5216=15216       P(X=3)=P(2).P(2).P(2)=16.16.16=1216Now,    E(X)=∑i=1npixi                              =0×125216+1×75216+2×15216+3×1216                              =0+75216+30216+3216=75+30+3216=108216=12Hence,  the  required    is  12.

Thumbs Up IconUpvote Thumbs Down Icon
New question posted•a year ago
New answer posted•a year ago
New answer posted•a year ago
New answer posted•a year ago
S
a year ago

This is a  Short Answer Type Questions as classified in NCERT Exemplar

Sol:

( i )     W e     k n o w     t h a t     P ( 0 ) + P ( 1 ) + P ( 2 ) + P ( 3 ) = 1 ⇒                                     k + k 2 + k 4 + k 8 = 1 ⇒                         8 k + 4 k + 2 k + k 8 = 1         ⇒ 1 5 k = 8 ∴                               k = 8 1 5 H e r e ,     n = 8 ,     p = 1 1 0 ,     q = 1 − 1 1 0 = 9 1 0 ( i i )     P ( X ≤ 2 ) = P ( X = 0 ) + P ( X = 1 ) + P ( X = 2 )                                                                 = k + k 2 + k 4 = 7 k 4 = 7 4 × 8 1 5 = 1 4 1 5 a n d     P ( X > 2 ) = P ( X = 3 ) = k 8 = 1 8 × 8 1 5 = 1 1 5 ( i i i )     P ( X ≤ 2 ) + P ( X > 2 ) = 1 4 1 5 + 1 1 5 = 1 5 1 5 = 1 .

...Read more

Thumbs Up IconUpvote Thumbs Down Icon
New answer posted•a year ago
New answer posted•a year ago
New answer posted•a year ago
S
a year ago

This is a  Short Answer Type Questions as classified in NCERT Exemplar

Sol:

H e r e ,     n = 7 ,     p = 0 . 2 5 = 2 5 1 0 0 = 1 4 ,     q = 1 − 1 4 = 3 4 ∴     P ( X ≥ 2 ) = 1 − [ P ( x = 0 ) + P ( x = 1 ) ]                                                     = 1 − [ C 0 7 ( 1 4 ) 0 ( 3 4 ) 7 + C 1 7 ( 1 4 ) 1 ( 3 4 ) 6 ]                                                     = 1 − [ ( 3 4 ) 7 + 7 4 ( 3 4 ) 6 ] = 1 − ( 3 4 ) 6 [ 3 4 + 7 4 ]                                                     = 1 − ( 3 4 ) 6 ( 1 0 4 ) = 1 − 7 2 9 4 0 9 6 × 1 0 4 = 1 − 7 2 9 0 1 6 3 8 4                                                     = 1 6 3 8 4 − 7 2 9 0 1 6 3 8 4 = 9 0 9 4 1 6 3 8 4 = 4 5 4 7 8 1 9 2 H e n c e ,     t h e     r e q u i r e d     p r o b a b i l i t y     i s     4 5 4 7 8 1 9 2 .

Thumbs Up IconUpvote Thumbs Down Icon
New answer posted•a year ago
S
a year ago

This is a  Short Answer Type Questions as classified in NCERT Exemplar

Sol:

Thumbs Up IconUpvote Thumbs Down Icon
New answer posted•a year ago
S
a year ago

This is a  Short Answer Type Questions as classified in NCERT Exemplar

Sol:

Here,    p=16+16+16=12⇒           q=1−12=12  and  n=5∴  P(x=r)=Crnprqn−r                         =C35(12)3(12)5−3=5!3!2!.(12)3.(12)2=10.18.14=516Hence,  the  required  probability  is  516.

Thumbs Up IconUpvote Thumbs Down Icon